OpenStudy (anonymous):

solve the system of equations 2x=4x-16 5y=4x+8

OpenStudy (anonymous):

-2x = 16 x = 8 5y = 4(8) + 8 5y = 40 y = 40/5 y = 8

OpenStudy (anonymous):

the top one which was my mistake is 2y=4x-16

OpenStudy (anonymous):

2y = 4x - 16 y = (4x-16)/2 y = 2x - 8 5 (2x -8) = 4x + 8 10x - 40 = 4x + 8 6x = 48 x = 48/6 x = 8 y = 2 (8) - 8 y = 8

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

you are welcome!

OpenStudy (anonymous):

can you help me with this one.? 9x-3y=36 7x-3y=45